OpenStudy (anonymous):

2cosx-2secx=3 Solve for 0 less than or equal to x less than 2pi

OpenStudy (anonymous):

\[\cos(x)-\frac{1}{\cos(x)}=\frac{3}{2}\] is a start

OpenStudy (anonymous):

\[2cosx - 2secx = 3 Solve for 0\le x < 2pi\]

OpenStudy (anonymous):

cosx=t \[2t^2-3t-2=0\]

OpenStudy (anonymous):

where did you get the squared from?

OpenStudy (anonymous):

\[(t-2)*(2t+1)=0\]

OpenStudy (anonymous):

\[t=2\quad and \quad t=-1/2\]

OpenStudy (anonymous):

there is no cosx^2 in this problem, and then you cannot factor it out

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

it is magic

OpenStudy (anonymous):

haha what does that mean then?

OpenStudy (anonymous):

trick is to put say \(\cos(x)=t\) as above and get \[t-\frac{1}{t}=\frac{3}{2}\] \[\frac{t^2-1}{t}=\frac{3}{2}\] \[2(t^2-1)=3t\] \[2t^2-2=3t\] \[2t^2-3t-2=0\] \[(t-2)(2t+1)=0\] \[t=2,t=-\frac{1}{2}\]

OpenStudy (anonymous):

then replace t by \(\cos(x)\)and get \(\cos(x)=2\) which is not possible or \(\cos(x)=-\frac{1}{2}\) which is possible

OpenStudy (anonymous):

on the interval \((0,2\pi)\) we have \[\cos(x)=-\frac{1}{2}\implies x=\frac{2\pi}{3}, x=\frac{4\pi}{3}\]

OpenStudy (anonymous):

Thankyou! that was a great help